OpenStudy (anonymous):

find the quotient: 4x^2-y^2/2x+y

OpenStudy (anonymous):

note that 4x^2-y^2 = (2x)^2 - y^2 = (2x+y)(2x-y)

OpenStudy (anonymous):

\[4x ^{2}-y ^{2}\div2x+4\]

OpenStudy (anonymous):

right, did you understand what I posted?

OpenStudy (anonymous):

\[4x ^{2}-y ^{2}\div2x+4\]

OpenStudy (anonymous):

not really

OpenStudy (anonymous):

do you understand that 4x^2 = (2x)^2 ?

OpenStudy (anonymous):

\[4x ^{2}-y ^{2}\div2x+4\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

okay. so, 4x^2 -y^2 = (2x)^2 - y^2 do you understand that?

OpenStudy (anonymous):

\[4x ^{2}-y ^{2}\div2x+4\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

okay, do you know the formula for \[a^2 - b^2\]

OpenStudy (anonymous):

\[4x ^{2}-y ^{2}\div2x+4\]

OpenStudy (anonymous):

i think so

OpenStudy (anonymous):

okay, what is it?